OpenStudy (anonymous):

Another precalc vector problem... Please help? If v=(1,2), w=(-1,2), and u=(5,12), for what scalar k will the vector kv+w be perpendicular to u?

OpenStudy (anonymous):

@ganeshie8 maybe you could help me with this one too? :)

ganeshie8 (ganeshie8):

start by finding the vector : kv+w

ganeshie8 (ganeshie8):

\(\large k\vec{v} + \vec{w} = k<1, 2> + <-1, 2>\) \(\large ~~~~~~~~~~= <k, 2k> + <-1, 2>\) \(\large ~~~~~~~~~~= <k-1, 2k+2>\)

ganeshie8 (ganeshie8):

fine ?

ganeshie8 (ganeshie8):

next use this : two vectors are perpendicular => dot product = 0

ganeshie8 (ganeshie8):

take the dot product : (kv+w) . (u) and set it equal to 0

OpenStudy (anonymous):

Oh I got! 5k - 5 + 24k +24 = 0 29k + 19 = 0 29k = -19 k = -19/29 So it's -19/29, which is the answer my book says. Wasn't nearly as hard as it seemed! Thanks again.